# A Context for Manifold Calculus

### v2

@@ 4. Classification of Homogeneous Functors
Let $M$ be a smooth manifold and let $I_{\mathrm{sm},M}^{=k}$ denote the subcategory of $\operatorname{Disj}_{\mathrm{sm}}^{\leq k}\left(M\right)$ spanned by the smooth isotopy equivalences between objects with exactly $k$ components. Corollary 4.5 remains valid if we replace $I_{M}^{=k}$ and by $I_{\mathrm{sm},M}^{=k}$ and $\operatorname{Homog}_{\mathrm{istp},\mathrm{exh}}^{k}\left(M;\mathcal{C}\right)$ by the full subcategory of $\operatorname{Fun}\left(N\left(\operatorname{Open}\left(M\right)\right),\mathcal{C}\right)$ spanned by the functors that are homogeneous of degree $k$, exhaustive, and carries smooth isotopy equivalences to equivalences, with essentially the same proof.
@@ 5. Context-Free Case
[…]
@@
In light of the constructions in non-context-free manifold calculus, some readers find it more natural to work with functors on $\mathsf{Mfld}_{n}$ which carry isotopy equivalences to equivalences, rather than functors defined on $\mathcal{M}\mathsf{fld}_{n}$. There is a reason for this: The localization theorem for isotopy equivalences (Theorem 2.24), which formed a basis of our argument in non-context-free manifold calculus, is significantly more subtle in the context-free case. We will come back to this point in Appendix A.
The following theorems, which are the main result of this section, are analogs of Theorems 3.4, 3.10 and 3.11 in the context-free case.
@@
    For each $n$-manifold $M$, each small $\infty$-category $\mathcal{I}$, and each functor $\chi:\mathcal{I}\to N\left(\operatorname{Open}\left(M\right)\right)$ satisfying the Weiss condition (Definition 3.9), the map
@@
Theorems 5.3 and 5.4 remain valid if we replace $\mathcal{M}\mathsf{fld}_{n}$, $\mathcal{D}\mathsf{isk}_{n}$, and $\mathcal{D}\mathsf{isk}_{n}^{\leq k}$ by $\mathcal{M}\mathsf{fld}_{\mathrm{sm},n}$, $\mathcal{D}\mathsf{isk}_{\mathrm{sm},n}$, and $\mathcal{D}\mathsf{isk}_{\mathrm{sm},n}^{\leq k}$, with essentially the same proof.
The proofs of Theorems 5.3 and 5.4 are very similar, so we will focus on Theorem 5.3.
@@
Clearly (2$\implies$(3$\implies$(4), so it suffices to show that (1$\implies$(2) and that (4$\implies$(1).
First we show that (1$\implies$(2). Suppose that $F$ is a left Kan extension of $F|\mathcal{D}\mathsf{isk}_{n}^{\leq k}$. Let $M$ be an $n$-manifold and let $\chi:\mathcal{I}\to\operatorname{Open}\left(M\right)$ be a functor of small categories satisfying the Weiss $k$-condition. We wish to show that the map
@@
is a left Kan extension of $N\left(\operatorname{Disj}^{\leq k}\left(M\right)\right)$. Since $F$ is a left Kan extension of $F|\mathcal{D}\mathsf{isk}_{n}^{\leq k}$, it suffices to show that, for each open set $U\subset M$, the map
$N\left(\operatorname{Disj}^{\leq k}\left(M\right)\right)_{/U}\to\mathcal{D}\mathsf{isk}_{n/U}^{\leq k}$
is final. This follows from Lemma 3.14 (or from Theorem 2.24, for localization functors are final [Lur24, Tag 02N9]).
Next we show that (4$\implies$(1). Let $G$ denote a left Kan extension of $F|\mathcal{D}\mathsf{isk}_{n}^{\leq k}$. We wish to show that the induced natural transformation $\alpha:G\to F$ is a natural equivalence. It suffices to show that, for each $n$-manifold $M$, the map $\alpha\iota:G\iota\to F\iota$ is a natural equivalence, where $\iota:N\left(\operatorname{Open}\left(M\right)\right)\to\mathcal{M}\mathsf{fld}_{n}$ denotes the inclusion. By Theorem 3.10, the functor $F\iota$ is a left Kan extension of its restriction to $N\left(\operatorname{Disj}^{\leq k}\left(M\right)\right)$. As we saw in the previous paragraph, the functor $G\iota$ is also left Kan extensions of its restriction to $N\left(\operatorname{Disj}^{\leq k}\left(M\right)\right)$. Since the components of $\alpha\iota$ at each object in $N\left(\operatorname{Disj}^{\leq k}\left(M\right)\right)$ are equivalences, we deduce that $\alpha\iota$ must be an equivalence. The proof is now complete. ∎
@@
Theorem 5.3 admits a following variant: Suppose we are given a right fibration $\pi:\mathcal{M}\to\mathcal{M}\mathsf{fld}_{n}$. Set $\mathcal{D}=\mathcal{D}\mathsf{isk}_{n}\times_{\mathcal{M}\mathsf{fld}_{n}}\mathcal{M}$ and $\mathcal{D}^{\leq k}=\mathcal{D}\mathsf{isk}_{n}^{\leq k}\times_{\mathcal{M}\mathsf{fld}_{n}}\mathcal{M}$. Given an $\infty$-category $\mathcal{C}$ with small colimits, we say that a functor $F:\mathcal{M}\to\mathcal{C}$ is a Weiss $k$-cosheaf if for each $n$-manifold $M$ and each functor $N\left(\operatorname{Open}\left(M\right)\right)\to\mathcal{M}$ rendering the diagram
${{\mathcal{M}}}$${{N(\operatorname{Open}(M))}}$${{\mathcal{M}\mathsf{fld}_{n}}}$
[…]

### v3

@@ 3. Classification of Homogeneous Functors
Notation 3.11 is motivated by the following observation: In the situation of Notation 3.11, suppose that $\mathcal{C}=\mathcal{S}$. By the straightening–unstraightening equivalence, $F$ corresponds to a Kan fibration $p:E\to\operatorname{Sing}X$. The naturality of the straightening–unstraightening correspondence gives an equivalence between $\Gamma\left(F;A\right)$ and the Kan complex $\operatorname{Fun}_{\operatorname{Sing}X}\left(\operatorname{Sing}A,E\right)$ of sections of $p$ over $\operatorname{Sing}A$.
We can also interpret $\operatorname{lim}_{\operatorname{Sing}X}F$ as a “twisted cochain.” To see this, consider the functors
$\mathbf{\Delta}\xleftarrow{\pi}\mathbf{\Delta}_{/\operatorname{Sing}X}\xrightarrow{\mathrm{fin}}\operatorname{Sing}X,$
where $\mathrm{fin}$ denotes the final vertex functor (dual of [Lan21, Lemma 3.3.6]). The functor $\mathrm{fin}$ is initial [Lan21, Theorem 3.3.8], so we have equivalences
$\operatorname{lim}_{\operatorname{Sing}X}F$ $\simeq\operatorname{lim}_{\mathbf{\Delta}_{/\operatorname{Sing}X}}\left(F\circ\mathrm{fin}\right)$
$\simeq\operatorname{lim}_{\mathbf{\Delta}}\operatorname{Ran}_{\pi}\left(F\circ\mathrm{fin}\right)$
$\simeq\operatorname{lim}_{\mathbf{\Delta}}\left([n]\mapsto\prod_{\sigma:\Delta^{n}\to X\in\left(\operatorname{Sing}X\right)_{n}}F\left(\sigma\left(n\right)\right)\right).$
Here the last equivalence is an instance of Lemma B.5. When $\mathcal{C}$ is the $\infty$-category of cochain complexes of abelian groups, the limit of a cosimplicial cochain complex is computed as the totalization of the associated bicomplex [Ara23b, Remark 1.20, Example 2.4]. If further $F$ takes values in the category of abelian groups, the resulting complex is nothing but the singular complex of $X$ with local coefficients $F$ [Whi12, VI.2].
Definition 3.13.
Let $M$ be a manifold, and let $k\geq 0$. We define the $k$-fold fat diagonal of $M$ by
$\blacktriangle_{k}\left(M\right)=\operatorname{SP}^{k}\left(M\right)\setminus B_{k}\left(M\right),$
where $\operatorname{SP}^{k}\left(M\right)=X^{k}/\Sigma_{k}$ denotes the $k$th symmetric product of $X$. We write $\operatorname{Nbd}\left(\blacktriangle_{k}\left(M\right)\right)$ for the poset of neighborhoods of $\blacktriangle_{k}\left(M\right)$ in $\operatorname{SP}^{k}\left(M\right)$.
If $M$ is triangulated and its vertices are well-ordered, $\operatorname{SP}^{k}\left(M\right)$ has a preferred triangulation containing $\blacktriangle_{k}\left(M\right)$ as its subcomplex (Example C.12). In this situation, we write $\operatorname{Nbd}^{\prime}\left(\blacktriangle_{k}\left(M\right)\right)$ for the full subposet of $\operatorname{Nbd}\left(\blacktriangle_{k}\left(M\right)\right)$ consisting of the (underlying polyhedra of) derived neighborhoods of $\blacktriangle_{k}\left(M\right)$ in $\operatorname{SP}^{k}\left(M\right)$.
Definition 3.14.
Let $\mathcal{C}$ be a pointed $\infty$-category with small limits and filtered colimits, let $M$ be a manifold that admits a PL structure, and let $F:\operatorname{Sing}B_{k}\left(M\right)\to\mathcal{C}$ be a functor. We define a functor $\Gamma^{\prime}_{\blacktriangle}\left(F;B_{k}\left(-\right)\right):\operatorname{Open}\left(M\right)^{\mathrm{op}}\to\mathcal{C}$ by
$\Gamma^{\prime}_{\blacktriangle}\left(F;B_{k}\left(U\right)\right)=\operatorname{colim}_{Q\in\operatorname{Nbd}\left(\blacktriangle_{k}\left(U\right)\right)^{\mathrm{op}}}\Gamma\left(F;Q\cap B_{k}\left(U\right)\right).$
Since $\operatorname{Nbd}\left(\blacktriangle_{k}\left(U\right)\right)^{\mathrm{op}}$ is weakly contractible (as it is filtered), there is a natural transformation
$\Gamma\left(F;B_{k}\left(-\right)\right)\simeq\operatorname{colim}_{Q\in\operatorname{Nbd}\left(\blacktriangle_{k}\left(-\right)\right)^{\mathrm{op}}}\Gamma\left(F;Q\cap B_{k}\left(-\right)\right)\to\Gamma^{\prime}_{\blacktriangle}\left(F;B_{k}\left(-\right)\right)$
We write $\Gamma_{\blacktriangle}\left(F;B_{k}\left(-\right)\right)$ for the fiber of this map.
Remark 3.15.
In the situation of Definition 3.14, the colimit defining $\Gamma^{\prime}_{\blacktriangle}\left(F;B_{k}\left(-\right)\right)$ exists. To see this, fix a triangulation of $U$. Since the inclusion $\operatorname{Nbd}^{\prime}\left(\blacktriangle_{k}\left(U\right)\right)\to\operatorname{Nbd}\left(\blacktriangle_{k}\left(U\right)\right)$ is initial (Lemma C.18), we only need to take the colimit over $\operatorname{Nbd}^{\prime}\left(\blacktriangle_{k}\left(U\right)\right)$. Since $\operatorname{Nbd}^{\prime}\left(\blacktriangle_{k}\left(U\right)\right)$ is weakly contractible (being cofiltered by Lemma C.18) and every inclusion $Q\to Q^{\prime}$ in $\operatorname{Nbd}^{\prime}\left(\blacktriangle_{k}\left(U\right)\right)$ induces a homotopy equivalence $Q\cap B_{k}\left(U\right)\xrightarrow{\simeq}Q^{\prime}\cap B_{k}\left(U\right)$, the colimit over $\operatorname{Nbd}^{\prime}\left(\blacktriangle_{k}\left(U\right)\right)$ indeed exists. Moreover, this argument shows that the map
$\Gamma\left(F;Q\cap B_{k}\left(U\right)\right)\to\Gamma^{\prime}_{\blacktriangle}\left(F;B_{k}\left(U\right)\right)$
is an equivalence for every $U\in\operatorname{Open}\left(M\right)$ and $Q\in\operatorname{Nbd}^{\prime}\left(\blacktriangle_{k}\left(U\right)\right)$.
We can now state the explicit formula for the inverse equivalence. In the case where $\mathcal{C}$ is the $\infty$-category of pointed spaces, the formula is due to Weiss, and the proof we present below should be compared with his original argument [Wei99, Section 7].
Theorem 3.16.
Let $\mathcal{C}$ be an $\infty$-category with small limits. For every manifold $M$ admitting a PL structure and every $k\geq 0$, the functor
$\Gamma_{\blacktriangle}\left(-;B_{k}\left(-\right)\right):\operatorname{Fun}\left(\operatorname{Sing}B_{k}\left(M\right),\mathcal{C}\right)\to\operatorname{Homog}_{\mathrm{istp},\mathrm{exh}}^{k}\left(M^{\mathrm{op}};\mathcal{C}\right)$
is well-defined and an inverse of the equivalence of Proposition 3.2.
Proof.
In light of Proposition 3.7, it will suffice to show that for each functor $F:\operatorname{Sing}B_{k}\left(M\right)\to\mathcal{C}$, the map $\Gamma\left(F;B_{k}\left(-\right)\right)\to\Gamma^{\prime}_{\blacktriangle}\left(F;B_{k}\left(-\right)\right)$ exhibits $\Gamma^{\prime}_{\blacktriangle}\left(F;B_{k}\left(-\right)\right)$ as a $\left(k-1\right)$th polynomial approximation of $\Gamma\left(F;B_{k}\left(-\right)\right)$. We will prove this in several steps. For notational convenience, we will write $\Phi=\Gamma\left(F;B_{k}\left(-\right)\right)$ and $\Psi=\Gamma^{\prime}_{\blacktriangle}\left(F;B_{k}\left(-\right)\right)$. Note that these functors can take closed subsets of $M$ as its inputs by extending them by the same defining formula.
  1. (Step1)
    We show that, for each $U\in\operatorname{Disj}\left(M\right)$, the map
    $\theta:\Gamma\left(F;B^{\prime\prime}_{k}\left(U\right)\right)\to\operatorname{colim}_{Q\in\operatorname{Nbd}\left(\blacktriangle^{k}\left(U\right)\right)}\Gamma\left(F;B_{k}\left(U\right)\cap Q\right)$
    is an equivalence. Identify each component of $U$ with $\mathbb{R}^{n}$ by choosing a homeomorphism. Given a decreasing continuous map $\varepsilon:[0,\infty)\to(0,1)$, let $\widetilde{Q}_{\varepsilon}\subset U^{k}$ denote the open set of the points $\left(p_{1},\dots,p_{k}\right)\in U^{k}$ satisfying the following conditions:
    • The map $\{p_{1},\dots,p_{k}\}\to\pi_{0}\left(U\right)$ is not injective.
    • For each component $U^{\prime}\subset U$ containing at least one point in $\{p_{1},\dots,p_{k}\}$, we have $\max_{p_{i},p_{j}\in U^{\prime}}\left|p_{i}-p_{j}\right|<\varepsilon\left(\max_{p_{i}\in U^{\prime}}\left|p_{i}\right|\right)$.
    Let $Q_{\varepsilon}\subset\operatorname{SP}^{k}\left(U\right)$ denote the image of $\widetilde{Q}_{\varepsilon}$. Notice that $Q_{\varepsilon}$ is a neighborhood of $\blacktriangle^{k}\left(U\right)$ in $\operatorname{SP}^{k}\left(U\right)$, and moreover that every neighborhood of $\blacktriangle^{k}\left(U\right)$ contains a neighborhood of the form $Q_{\varepsilon}$. Thus we obtain an equivalence
    $\operatorname{colim}_{\varepsilon}\Gamma\left(F;Q_{\varepsilon}\cap B_{k}\left(U\right)\right)\xrightarrow{\simeq}\operatorname{colim}_{Q\in\operatorname{Nbd}\left(\blacktriangle^{k}\left(U\right)\right)}\Gamma\left(F;Q\cap B_{k}\left(U\right)\right).$
    On the other hand, the inclusion $Q_{\varepsilon}\cap B_{k}\left(U\right)\to B^{\prime\prime}_{k}\left(U\right)$ is a homotopy equivalence. To see this, choose a homeomorphism $f:[0,\infty)\to[0,1)$. By using, for each component $U^{\prime}\subset U$ intersecting $\{p_{1},\dots,p_{k}\}$, the map
    $\{p_{i}\}_{i\in U^{\prime\prime}}\mapsto\{f\left(\max_{p_{s}\in U^{\prime}}\left|p_{s}\right|\right)\varepsilon\left(\max_{p_{s}\in U^{\prime}}\left|p_{s}\right|\right)p_{i}\}_{i\in U^{\prime\prime}},$
    we obtain a map $B^{\prime\prime}_{k}\left(U\right)\to Q_{\varepsilon}\cap B_{k}\left(U\right)$. Using linear homotopies, one verifies that this is an inverse homotopy equivalence of the inclusion. (The above map and the homotopies are well-defined because $\varepsilon$ is a decreasing function with values in $\left(0,1\right)$.) So we obtain another equivalence
    $\Gamma\left(F;B^{\prime\prime}_{k}\left(U\right)\right)\xrightarrow{\simeq}\operatorname{colim}_{\varepsilon}\Gamma\left(F;Q_{\varepsilon}\cap B_{k}\left(U\right)\right).$
    The map $\theta$ is the composite of these maps, so it is an equivalence, as required.
  2. (Step2)
    Let $T_{k-1}\Phi$ and $T_{k-1}\Psi$ denote the right Kan extensions of $\Phi$ and $\Psi$ along the inclusion $\operatorname{Disj}^{\leq k-1}\left(M\right)^{\mathrm{op}}\hookrightarrow\operatorname{Open}\left(M\right)^{\mathrm{op}}$. We then have the following diagram:
    $\Phi$$\Psi$$T_{k-1}\Phi$$T_{k-1}\Psi$$\scriptstyle{\eta_{\Phi}}$$\scriptstyle{\eta_{\Psi}}$$\scriptstyle{\alpha}$
    The map $\alpha$ is an equivalence by Step 1 (because $B^{\prime\prime}_{k}\left(U\right)=B_{k}\left(U\right)$ for every $U\in\operatorname{Disj}^{\leq k-1}\left(M\right)$). Consequently, it will suffice to show that $\eta_{\Psi}$ is an equivalence. We will prove this in the next step.
  3. (Step3)
    We show that $\eta_{\Psi}$ is an equivalence. By Remark 1.15, it suffices to show that every open set $U\subset M$ equipped with a triangulation coming from a PL structure of $M$ satisfies the following conditions:
    1. (I)
      If $U\in\operatorname{Disj}\left(M\right)$, then the map $\eta_{\Psi,U}:\Psi\left(U\right)\to T_{k-1}\Psi\left(U\right)$ is an equivalence.
    2. (II)
      Let $K_{0}\subset K_{1}\subset\cdots\subset U$ be a sequence of subcomplexes whose interiors cover $U$. The map
      $\Psi\left(U\right)\to\operatorname{lim}_{i}\Psi\left(K_{i}\right)$
      is an equivalence.
    3. (III)
      Let $A_{0},\dots,A_{k-1}\subset U$ be pairwise disjoint compact subcomplexes. The map
      $\Psi\left(U\right)\to\operatorname{lim}_{S\in\mathcal{P}_{0}\left(\{0,\dots,k-1\}\right)}\Psi\left(U\setminus\bigcup_{i\in S}A_{i}\right)$
      is an equivalence.
    For (I), we consider the diagram
    $\Phi(U)$$\Gamma(F;B^{\prime\prime}_{k}(U))$$\Psi(U)$$\operatorname{lim}_{V\in\operatorname{Disj}^{\leq k-1}(U)}\Gamma(F;B_{k}(V))$$T_{k-1}\Psi(U).$$\scriptstyle{\eta_{\Phi,U}}\scriptstyle{\simeq}$$\scriptstyle{\psi}\scriptstyle{\simeq}$$\scriptstyle{\phi}$$\scriptstyle{\eta_{\Psi,U}}$$\scriptstyle{\alpha_{U}}\scriptstyle{\simeq}$
    We saw in the proof of Proposition 3.7 that $\phi$ is an equivalence, and in Step 1 that $\psi$ is an equivalence. Thus $\eta_{\Psi,U}$ is an equivalence, as claimed.
    For (II), use Example C.12 and triangulate $\operatorname{SP}^{k}\left(U\right)$ in such a way that it contains $\blacktriangle_{k}\left(U\right)$ as a subcomplex. As we observed in Remark 3.15, for an arbitrary element $Q\in\operatorname{Nbd}^{\prime}\left(\blacktriangle_{k}\left(U\right)\right)$, the map
    $\Gamma\left(F;Q\cap B_{k}\left(U\right)\right)\to\Psi\left(U\right)$
    is an equivalence. A similar argument shows that the map
    $\Gamma\left(F;Q\cap B_{k}\left(K_{i}\right)\right)\to\Psi\left(K_{i}\right)$
    is an equivalence for every $i$, since $Q\cap\operatorname{SP}^{k}\left(K_{i}\right)\in\operatorname{Nbd}^{\prime}\left(\blacktriangle_{k}\left(K_{i}\right)\right)$ (Remark C.17). Consequently, we are reduced to showing that the map
    $\Gamma\left(F;Q\cap B_{k}\left(U\right)\right)\to\operatorname{lim}_{i}\Gamma\left(F;Q\cap B_{k}\left(K_{i}\right)\right)$
    is an equivalence. For this, it suffices to show that $\operatorname{Sing}\left(Q\cap B_{k}\left(U\right)\right)=\bigcup_{i\geq 0}\operatorname{Sing}\left(Q\cap B_{k}\left(K_{i}\right)\right)$ (by Remark B.10 and the fact filtered colimits of simplicial sets already compute homotopy colimits), which follows by inspection.
    For (III), subdividing $U$ if necessary, we may assume that no simplex of $U$ intersects $A_{i}$ and $A_{j}$ for distinct indices $i\neq j$. Let $C_{i}$ denote the simplicial complement of $A_{i}$ in $U$. For each subset $S\subset\{0,\dots,k\}$, the inclusion $U\setminus\bigcup_{i\in S}A_{i}\to C_{i}$ is a homotopy equivalence (see Remark C.15), so it suffices to show that the map
    $\Psi\left(U\right)\to\operatorname{lim}_{S\in\mathcal{P}_{0}\left(\{0,\dots,k-1\}\right)}\Psi\left(C_{i}\right)$
    is an equivalence. This follows from an argument similar to the one in the previous paragraph (and [Lur17, Theorem A.3.1]), noting that if $Q\in\operatorname{Nbd}\left(\blacktriangle_{k}\left(U\right)\right)$ is sufficiently small, we have $B_{k}\left(U\right)\cap Q=\bigcup_{i=0}^{k-1}B_{k}\left(C_{i}\right)\cap Q$. ∎
As a corollary of Proposition 3.7, we give a connectivity estimate of polynomial approximations. When $\mathcal{C}$ is the category of chain complexes of abelian groups or that of spectra, this is proved in [RW14, Lemma E.4] by a rather ad-hoc argument that reduces the claim to that of pointed spaces. Proposition 3.7 allows us to get away with this reduction while keeping the essence of the argument of [RW14] intact.
Corollary 3.17.
Let $M$ be an $n$-manifold admitting a PL structure, let $\mathcal{C}$ be a stable $\infty$-category with small limits and a t-structure, and let $F:\operatorname{Open}\left(M\right)^{\mathrm{op}}\to\mathcal{C}$ be an isotopy functor. Suppose that:
  1. (i)
    For each $U\in\operatorname{Disj}\left(M\right)$, the object $F\left(U\right)\in\mathcal{C}$ is $a$-connected.
  2. (ii)
    0-connected objects are stable under small products.
Then for every $k\geq 0$, the values of $T_{k}F$ are $\left(a-kn-k\right)$-connected.
Proof.
Shifting degrees if necessary, we may assume that $a=0$. We set $\lambda_{k}=-kn-k$. The proof proceeds by induction on $k$.
For the base step where $k=0$, $T_{0}F$ is the constant diagram at $F\left(\emptyset\right)$, so the claim is trivial. For the inductive step, let $k\geq 1$, and suppose we have proved the claim for $T_{k-1}F$. We have a fiber sequence
$L_{k}F\to T_{k}F\to T_{k-1}F$
in $\operatorname{Fun}\left(\operatorname{Open}\left(M\right)^{\mathrm{op}},\mathcal{C}\right)$. By the induction hypothesis, the values of $T_{k-1}F$ are $\lambda_{k-1}$-connected, and hence are $\lambda_{k}$-connected. Thus, by the long exact sequence of homotopy groups, we are reduced to showing that $L_{k}F$ is $\lambda_{k}$-connected. We prove this in several steps.
  1. (Step1)
    We show that, for each $U\in I_{M}^{=k}$, the object $\left(L_{k}F\right)\left(U\right)\in\mathcal{C}$ is $\left(-k\right)$-connected. By Remark 3.6 and condition (i), $\left(L_{k}F\right)\left(U\right)$ can be written as a total fiber of a diagram $1]^{k}\to\mathcal{C}_{\geq 1}$. Since the fiber of maps of $a$-connected objects is $\left(a-1\right)$-connected, the claim thus follows by induction on $k$ and the inductive formula for total fibers [ACB22, Corollary 2.2].
  2. (Step2)
    We will show that the values of $L_{k}F$ are $\lambda_{k}$-connected. Let $H:\operatorname{Sing}B_{k}\left(M\right)\to\mathcal{C}$ denote the functor corresponding to the homogeneous functor $L_{k}F$. For each $U\in\operatorname{Open}\left(M\right)$, Proposition 3.16 gives us an equivalence
    $L_{k}F\left(U\right)$ $\simeq\Gamma_{\blacktriangle}\left(H;U\right)$
    $\simeq\mathrm{fib}\left(\Gamma\left(H;B_{k}\left(U\right)\right)\to\Gamma\left(H;B_{k}\left(U\right)\cap Q\right)\right)$
    $\simeq\mathrm{fib}\left(\operatorname{lim}_{\operatorname{Sing}B_{k}\left(U\right)}H\to\operatorname{lim}_{\operatorname{Sing}\left(B_{k}\left(U\right)\cap Q\right)}H\right),$
    where $Q$ is a derived neighborhood of $\blacktriangle_{k}\left(U\right)$ in $\operatorname{SP}^{k}\left(U\right)$ with respect to some triangulation of $U$. Since $B_{k}\left(U\right)$ is obtained from $B_{k}\left(U\right)\cap Q$ by attaching cells of dimension $\leq kn$, combining Step 1 with Lemma B.17 (and condition (ii)), we deduce that the map
    $\operatorname{lim}_{\operatorname{Sing}B_{k}\left(U\right)}H\to\operatorname{lim}_{\operatorname{Sing}\left(B_{k}\left(U\right)\cap Q\right)}H$
    is $\left(\lambda_{k}+1\right)$-connected. Thus its fiber is $\lambda_{k}$-connected, completing the induction.∎
Remark 3.18.
[…]
@@ 4. Context-Free Case
    We say that $F$ is a Weiss $k$-cosheaf (resp. Weiss cosheaf) if the composite $\mathsf{Mfld}_{n}\to\mathcal{M}\mathsf{fld}_{n}\xrightarrow{F}\mathcal{C}$ is a cosheaf with respect to the Weiss $k$-topology (resp. Weiss topology).
@@
In light of the constructions in non-context-free manifold calculus, some readers find it more natural to work with functors on $\mathsf{Mfld}_{n}$ which carry isotopy equivalences to equivalences, rather than functors defined on $\mathcal{M}\mathsf{fld}_{n}$. There is a reason for this: The localization theorem for isotopy equivalences (Theorem A.1), which formed a basis of our argument in non-context-free manifold calculus, is significantly more subtle in the context-free case. We will come back to this point in Appendix A.
Remark 4.3.
The Taylor cotower of $F$ coincides with the tower of Kurannich–Kupers [KK24, 5.3.3], both in the smooth case and the topological case. See [KK24, Remark 5.8].
The following theorems, which are the main result of this section, are analogs of Theorems 1.3, 1.9 and 1.10 in the context-free case.
@@
    For each $n$-manifold $M$, each small $\infty$-category $\mathcal{I}$, and each functor $\chi:\mathcal{I}\to\operatorname{Open}\left(M\right)$ satisfying the Weiss condition (Definition 1.8), the map
@@
Theorems 4.4 and 4.5 remain valid if we replace $\mathcal{M}\mathsf{fld}_{n}$, $\mathcal{D}\mathsf{isk}_{n}$, and $\mathcal{D}\mathsf{isk}_{n}^{\leq k}$ by $\mathcal{M}\mathsf{fld}_{\mathrm{sm},n}$, $\mathcal{D}\mathsf{isk}_{\mathrm{sm},n}$, and $\mathcal{D}\mathsf{isk}_{\mathrm{sm},n}^{\leq k}$, with essentially the same proof.
The proofs of Theorems 4.4 and 4.5 are very similar, so we will focus on Theorem 4.4.
@@
[…]
@@
$\operatorname{Open}\left(M\right)\to\mathcal{M}\mathsf{fld}_{n}\xrightarrow{F}\mathcal{C}$
is a left Kan extension of $\operatorname{Disj}^{\leq k}\left(M\right)$. Since $F$ is a left Kan extension of $F|\mathcal{D}\mathsf{isk}_{n}^{\leq k}$, it suffices to show that, for each open set $U\subset M$, the map
$\operatorname{Disj}^{\leq k}\left(M\right)_{/U}\to\mathcal{D}\mathsf{isk}_{n/U}^{\leq k}$
is final. This follows from Lemma 1.11 (or from Theorem A.1, for localization functors are final [Lur25, Tag 02N9]).
@@
Theorem 4.4 admits a following variant: Suppose we are given a right fibration $\pi:\mathcal{M}\to\mathcal{M}\mathsf{fld}_{n}$. Set $\mathcal{D}=\mathcal{D}\mathsf{isk}_{n}\times_{\mathcal{M}\mathsf{fld}_{n}}\mathcal{M}$ and $\mathcal{D}^{\leq k}=\mathcal{D}\mathsf{isk}_{n}^{\leq k}\times_{\mathcal{M}\mathsf{fld}_{n}}\mathcal{M}$. Given an $\infty$-category $\mathcal{C}$ with small colimits, we say that a functor $F:\mathcal{M}\to\mathcal{C}$ is a Weiss $k$-cosheaf if for each $n$-manifold $M$ and each functor $\operatorname{Open}\left(M\right)\to\mathcal{M}$ rendering the diagram
$\mathcal{M}$$\operatorname{Open}(M)$$\mathcal{M}\mathsf{fld}_{n}$}}
commutative, the composite $\operatorname{Open}\left(M\right)\to\mathcal{M}\xrightarrow{F}\mathcal{C}$ is a Weiss $k$-cosheaf. We define what it means for a functor $\mathcal{M}\to\mathcal{C}$ to be a Weiss cosheaf, $k$-excisive, or exhaustive similarly. Then the following conditions for a functor $F:\mathcal{M}\to\mathcal{C}$ are equivalent: